a square rug has an area of 91 square feet.
what is a good approximation of the perimeter of the rug?

Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

if the rug is square, then the formula for the area of the rug is A = S^2.

since A = 91, then the formula becomes 91 = S^2.

solve for S to get S = sqrt(91).

the formula for the perimter of the square rug is 4S = P

since S = sqrt(91), then the formula becomes 4 * sqrt(91) = P.

the perimeter is equal to 4 * sqrt(91).

sqrt(91) is approximately equal to 9.539392014 feet.

4 * sqrt(91) is approximately equal to 38.15756806 feet.


Answer by ikleyn(52778)

.
Good approximation is 38.16 feet.



38.15756806 feet is not a good approximation.